OpenStudy (hellokitty17):

@phi Pentagon PQRST and its reflection, pentagon P'Q'R'S'T', are shown in the coordinate plane below: What is the line of reflection between pentagons PQRST and P'Q'R'S'T'? x = 1 y = x x = 0 y = 0

OpenStudy (anonymous):

is there a coordinate plane to look at?

OpenStudy (phi):

what line is halfway between the two figures?

OpenStudy (hellokitty17):

y

OpenStudy (phi):

if you connect R to R' with a line, and find the halfway point. and do that for S to S'. put a dot halfway between both points will be "somewhere" what are their (x,y) values?

OpenStudy (hellokitty17):

the y axis

OpenStudy (phi):

do you know the equation of the y-axis?

OpenStudy (hellokitty17):

no sorry

OpenStudy (anonymous):

x=0

OpenStudy (phi):

you can figure it out. (0,0) and (1,0) are two points on the y-axis what is the slope?

OpenStudy (hellokitty17):

0

OpenStudy (phi):

slope is change in y divided by change in x (0-0)/(1-0) = 0/1 = 0 y = mx + b m is 0 y = 0*x + b y = b to find b plug in (0,0) there is no x, but it does not matter. y is 0 0= b so y= 0

OpenStudy (phi):

if you are "on the y-axis" the y value is always 0, and x can be any number

OpenStudy (hellokitty17):

thank you! :) I will tag you for the next question